OpenStudy (anonymous):

I need help what is the answer of this (x)+(17cm)=15cm

OpenStudy (anonymous):

Negative numbers can be confusing. Even though you're dealing with centimeters, try to think about it as temperature for a minute. If you're at 17 degrees, how do you get to 15 degrees? You go down 2. (Another way of saying that is to add -2). x + 17 = 15 x+ 17 -17 = 15 -17 x = -2 cm (That means you go 2 cm to the left.)

OpenStudy (anonymous):



OpenStudy (anonymous):

x=-2cm